OpenStudy (idealist10):

Evaluate the double integral ye^x dA, where R is the triangular region with vertices (0, 0), (4, 4), and (6, 0).

OpenStudy (idealist10):

@ganeshie8 @iambatman @wio @Loser66 @Destinymasha @paki @Compassionate @Kainui

ganeshie8 (ganeshie8):

As a start, sketch the region in xy plane

OpenStudy (loser66):



OpenStudy (loser66):

For this problem, you need change a little bit, the equation of the line OA is x =y the equation of the line AB is y = -2x+12 or \(x= 6-\dfrac{y}{2}\) and fixed y, limit from 0 to 4 Hence your integral is \[\int_0^4\int_{y}^{6-(y/2)} ye^xdxdy\]

OpenStudy (idealist10):

How did you get y=-2x+12?

OpenStudy (loser66):

find out the slope of the line by formula \(\dfrac{y-y_0}{x-x_0}\) then plug back to the equation of the line y = mx +b to get it

OpenStudy (loser66):

Don't get??

OpenStudy (loser66):

Ok, 2 points (4,4) , (6,0) , hence the slope \(m =\dfrac{y-y_0}{x-x_0}=\dfrac{0-4}{6-4}=-2\) pick one of the point to plug in the equation, I pick (6,0) , hence \( y = mx +b \\0= -2*6+b\\b= 12\\put ~~back\\y =-2x+12\\x=6-\dfrac{y}{2} \)

OpenStudy (loser66):

got to go. Good luck

OpenStudy (idealist10):

But how did you get x=y?

ganeshie8 (ganeshie8):